1) What is the magnitude of the final momentum of

1) What is the magnitude of the final momentum of the system (all three pieces)? 0

2) What is the mass of the third piece? 6.6

3) What is the x-component of the velocity of the third piece? 10.6

I NEED HELP IN THESE THREE PLEASE

4) What is the y-component of the velocity of the third piece?

5) What is the magnitude of the velocity of the center of mass of the pieces after the collision?

6) Calculate the increase in kinetic energy of the pieces during the explosion?
Answer #1

4) let y componenet be v

by conserving momentum in y direction we have

0= m1*v1sin26.7 -m2*v2*sin25 +m3*v

hence

v = 42.92+47.93 / 6.6

V = -13.76 m/sec

5)center of mass velocity after collision is = 0

because of conservation of momentum

6) increase in K.e = 1/2*(4.7*26.72 + 5.3*21.42 + 6.6*17.372) = 3884.619 joules